OpenStudy (anonymous):

Solve for v. m=mbv/k a. v=mb/k b. v=mkb c. v=b/mk d. v=mk/b Tue Dec 16 2014 12:41

OpenStudy (kamibug):

Are you sure you copied everything correctly? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

None of these is correct, the answer isn't dependent on m if the equation is m=mbv/k

OpenStudy (anonymous):

\[m = \frac{ mbv }{ k } \] multiply by k: \[mk=mbv\] divide by m: \[k=bv\] divide by b: \[v=\frac{ k }{ b }\]

OpenStudy (anonymous):

Even if you kept the m, it would still be: \[v = \frac{ mk }{ mb }\]

OpenStudy (kamibug):

The question is broken. :P

OpenStudy (anonymous):

no word by word and number by number